Rigo:

Mr. Hauseman has 17 students in his class, three of whom are freshmen, and the rest are from other classes. He is going to draw two students randomly to be partners.

justus:

Seems like there is more to your question, please post it so we can further help you :)

Rigo:

3/17 are freshmen If there are x juniors, then P(freshman,junior) = 3/17 * x/16 = 3x/272 = 9/136 so, x=6

justus:

Yes

Rigo:

sorry, i just got stuck thx

justus:

Np, is there anything I can help you with.

Rigo:

Sara is determining the total cost of her meal. The meal itself costs $9.75, and Sara would like to leave a 15% tip. She uses the expression $9.75 + $9.75(0.15) to determine the total cost. Which expression could Sara use to make the calculation easier?

Rigo:

whats an expression

justus:

$9.75(1.15)

Rigo:

thx

justus:

Yep

justus:

If you are done with this question, you can click the close button :)

isis:

whoa your a human calculator